LUCI/LBC Technical Observing, prep for Partner Science:
UT 2011 Feb 26

Observers: M.Pedani. A.Rakich,. O.Kuhn from downtown
Telescope Operator: S.Allanson
Instrument Support: T.Shih
Telescope Support: J.Urban
SW support: D.Cox

Summary

  • Taken 1 hour of Field Check Data Collection

22:55 Sky is mostly coudy with few chances to open. LUCI ready to go

01:20 Weather: Many clouds with some patches of clear sky. T=-3.2 RH=37% T_mirror=-3.5 Wind=12 m/s. Vent doors opened mid-way

02:19 Weather: Patches of clouds, T=-3.5 RH=34% T_mirror=-2.9 Wind=16m/s.

02:20 Check Pointing IE=-86.0 CA=70.0

02:34 Slew to BS9122 to get Field Aberration Data. Seeing is ~ 1.0"

02:38 field_check, /GCS on BS9122 (stars in: BS9122_short_field.list)

02:39 ===> Error in irc_presetTelescope. We restart GCS and re-try First Image: left_wfscimage000019.fits Last Image: left_wfscimage000020.fits

02:43 Script aborts again! We call Torsten.

02:53 Re-start script: field_check, /GCS on BS9122 First Image: left_wfscimage000021.fits Last Image: left_wfscimage000083.fits

03:56 Weather: T=-3.8 T_mirror=-2.9 RH=34% Wind=15-17 m/s

04:05 Slew to B3.2 Source

DJT 01Mar2011 - The pixel offsets of the dither sequence were compared to the requested 
DETXY offsets (in arcsec but scaled to pixels by dividing by 0.12) using iraf's GEOMAP task.  
The resulting pattern shows a mean drift of 1.2 pixels in X and 0.2 pixels in Y across the 
sequence, possibly from flexure.    Without correcting for this drift, the fit shows an 0.25 
degree rotation.  This suggests that the AGw XY coordinate system is misaligned with the 
DETXY system by this amount, but it could also be the LUCI LEFTZEROPOINT is off.  Linear 
fits to the X and Y sequence data suggest the rms positioning is <~40mas.

### GEOMAP output (from Xoff Yoff Xrequested Yrequested data):
Coordinate mapping status
    Xin and Yin fit rms: 0.6204616  0.2862498
Coordinate mapping parameters
    Mean Xref and Yref: 111.934  125.4052
    Mean Xin and Yin: -5.1164  6.9332
    X and Y shift: -116.6751  -119.2271  (xin  yin)
    X and Y scale: 1.001657  1.002049  (xin / xref  yin / yref)
    X and Y axis rotation: 359.74481  359.74481  (degrees  degrees)

05:00 Switch to LBC to tweak DOFPIA code (Andrew+Olga from Downtown)

05:05 Weather: T=-4.3 T_mirror=-3.2 RH=35% Wind=15m/s still some cirrus clouds

05:30 Switch complete. Turning on LBCB

05:39 Warmed up and then turned on LBCB. Slew to a field at 9h 55deg (elev=67 deg) 054053. Then remove global z offset of 0.9 and repeat, went too far and to other side.

05:44 Repeat with global z offset = 0.6mm 054615 Good.

05:46 Try now a global offset for Rx of -10".

05:50 Start dofpia: dofpia, /first, BlueLimit=10, /blueonly Limit is applied to z4,z5,z6,z11

Checking for Blue Convergence
Limits:     10.0    10.0    10.0   400.0   400.0    10.0    50.0

06:04 Slewing to another field with more stars - we had been fitting 2-4 pupils. Going to cluster, Palomar 4, Ursa Major Dwarf 062140 about 5 pupils found

06:22 Slew to field around BS9148, 6 pupils found. Letting dofpia continue...when near convergence, drive in aberrations:

06:34 when 063419 reading out, drive in z11=+1000

  292    306     93    184    281    -47      8    y  1.15  0.23                      /newdata/lbcb.20110226.063419.fits  V-BESSEL
drove in z11=+1000nm
   313   -233    -61    289    107   -703    -81    y  1.00  0.49                      /newdata/lbcb.20110226.063536.fits  V-BESSEL
  -762    175   -297   -362   -264    -43     50    y  1.14  0.38                      /newdata/lbcb.20110226.063654.fits  V-BESSEL
   302   -133    133    -58   -109   -205     23    y  1.14  0.21                      /newdata/lbcb.20110226.063809.fits  V-BESSEL
   943   -178    146    -64    -12   -329    -43    y  1.18  0.44                      /newdata/lbcb.20110226.063923.fits  V-BESSEL
  -637    282   -245     -4    -36     63     19    y  1.10  0.29                      /newdata/lbcb.20110226.064037.fits  V-BESSEL
  -810   -189    268     24     50    202     51    y  1.19  0.37                      /newdata/lbcb.20110226.064151.fits  V-BESSEL
    31      0      0      0      0    -16      0    y  0.00  0.00                     dofpia_sph3_adjustment       
Stopped with "S"

06:44 Changed -2.4 to +1.0 in BlueZIMatrix: restart dofpia: 064553 is first

06:48 Send z11=+1000 after 064940, then after 'converged' sent z11=+1000 after 065825

 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
   298    293    -54    192     15   -276    -31    y  1.16  0.26                      /newdata/lbcb.20110226.064553.fits  V-BESSEL
  -565   -238    258   -100    -75    208    100    y  1.45  0.29                      /newdata/lbcb.20110226.064709.fits  V-BESSEL
     2    184   -441    -56   -115    114     93    y  1.42  0.19                      /newdata/lbcb.20110226.064825.fits  V-BESSEL
  1039   -242    262      6    127   -544    -73    y  1.29  0.55                      /newdata/lbcb.20110226.064940.fits  V-BESSEL
   737    185   -457    178    -18  -1000   -100    y  1.13  0.72                      /newdata/lbcb.20110226.065056.fits  V-BESSEL
  -722     49    366     22     59   -117     16    y  1.24  0.32                      /newdata/lbcb.20110226.065215.fits  V-BESSEL
   847      6   -126     88     78   -198    -39    y  1.13  0.37                      /newdata/lbcb.20110226.065329.fits  V-BESSEL
   245   -439   -168    -67     37    -60    -30    y  1.24  0.19                      /newdata/lbcb.20110226.065442.fits  V-BESSEL
   -66    295     17     43    -15     10     -2    y  1.22  0.12                      /newdata/lbcb.20110226.065557.fits  V-BESSEL
     3     39    240   -293    -20    244     26    y  1.28  0.23                      /newdata/lbcb.20110226.065711.fits  V-BESSEL
  -844    373    -32    178    -36   -150      8    y  1.28  0.38                      /newdata/lbcb.20110226.065825.fits  V-BESSEL

07:05 Going to a new target, M3. Going S away from core. Set Z4/Z11 matrix element to +10.0 (ridiculous to make sure our changes are having some effect).

07:20 Make a new offset, 7' south, and run again with +10. 072529... a little further offset, 10' south.

07:32 dofpia Failed after 072650 with error message

Waiting for OB to finish and data files to appear in /newdata
Processing: /newdata/lbcb.20110226.073225.fits
% Compiled module: ERRORLOG.
% GETIMAGE: Program caused arithmetic error: Floating illegal operand
% Execution halted at: GETIMAGE           72
  /home/lbcobs/LBCFPIA/lbcfpia/src/getimage.pro
%                      LBCFPIA           215
  /home/lbcobs/LBCFPIA/lbcfpia/src/lbcfpia.pro
%                      DOFPIA            444
  /home/lbcobs/LBCFPIA/lbcfpia/src/dofpia.pro
%                      $MAIN$          

07:33 With Matrix element = +10.0 we see z4 change is still flipped in sign to z11 change, and more exaggerated than for -2.4. We will try going the other way, with the matrix element 12 < -2.4. Changed to -4 Added z11 = +1000 nm before 075131.

 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  4920   -315     21   -140   -123   -458    -31    y  1.07  1.97                      /newdata/lbcb.20110226.074612.fits  V-BESSEL
    84    116   -282    181    111   -190      1    y  0.99  0.20                      /newdata/lbcb.20110226.074735.fits  V-BESSEL
   291   -244    140    -80    -12   -204    -11    y  0.96  0.21                      /newdata/lbcb.20110226.074856.fits  V-BESSEL
  -149      6    259     51     27    -83    -11    y  1.02  0.14                      /newdata/lbcb.20110226.075015.fits  V-BESSEL
   785    314    101    111     70   -659    -70    y  0.83  0.54                      /newdata/lbcb.20110226.075131.fits  V-BESSEL
  -445   -583    -86     29   -105   -162     47    y  0.96  0.28                      /newdata/lbcb.20110226.075251.fits  V-BESSEL
   -72    204   -170    -65     76   -157     -3    y  0.95  0.16                      /newdata/lbcb.20110226.075413.fits  V-BESSEL
  -434    170    -88     44     -2     79     59    y  1.14  0.20                      /newdata/lbcb.20110226.075535.fits  V-BESSEL
   -93    228    247    -50     -5    -15     14    y  1.10  0.13                      /newdata/lbcb.20110226.075652.fits  V-BESSEL
S to stop

07:57 Try -5 now. Starting at 080215

08:08 080855 sent again the last zernikes before this one.

08:11 081132 after this z11=+1000 was sent. We are noticing some Z4/Z5 correlations, large values of same sign for both. We need to look at this more --- could it be windshake?

08:21 after 082212 sending -1000 Z11.

08:22 082321 is next image after Z11 bump. Seemed like -4 was better than -5... Going back to try -4 with giving Z11 = - 1000 nm

    24   -221    230   -189     65    345     66    y  1.45  0.27                      /newdata/lbcb.20110226.082321.fits  V-BESSEL
  -716    252   -313     97   -152     63    -80    y  1.08  0.33                      /newdata/lbcb.20110226.082439.fits  V-BESSEL
    65    253   -333     33     28    148    -44    y  1.13  0.18                      /newdata/lbcb.20110226.082600.fits  V-BESSEL
   179   -231    349   -114     13    223     47    y  1.26  0.22                      /newdata/lbcb.20110226.082720.fits  V-BESSEL
  1402    235   -166    125    -55   -328    -99    y  1.02  0.60                      /newdata/lbcb.20110226.082838.fits  V-BESSEL
  -482   -338    162     47    -34   -117    -62    y  0.94  0.25                      /newdata/lbcb.20110226.082959.fits  V-BESSEL
  -681    292   -271     12     49    217     27    y  1.04  0.33                      /newdata/lbcb.20110226.083120.fits  V-BESSEL
   305   -192    279    -86    -55     94     23    y  1.16  0.19                      /newdata/lbcb.20110226.083237.fits  V-BESSEL
   232    223   -116    -26     31    -70    -26    y  1.07  0.15                      /newdata/lbcb.20110226.083356.fits  V-BESSEL
  -113   -349    330     74     -5    101     47    y  1.22  0.18                      /newdata/lbcb.20110226.083515.fits  V-BESSEL
  -453    167   -261     11    -23     91     21    y  1.13  0.22                      /newdata/lbcb.20110226.083634.fits  V-BESSEL
    -7      0      0      0      0      4      0    y  0.00  0.00                     dofpia_sph3_adjustment             

5805.

08:37 Change to -4.0 and run dofpia again. 083924...

08:47 Sending +1000 nm Z11 before 084841. Converged to 250/400... Sending -1000nm Z11 after 085805. Slow to correct negative z11 (as we saw also when the matrix element was -5. But again the positive z11 corr are accompanied by negative z4, like when we went to +10.

09:02 Converging. Now change 12 element to -6.0. Not converging well... Stopped with "S".

09:16 Commenting out Z4Z11Factor = -1.9 in lbcfpia.pro and changing lbcfpia.cfg file value of Z4Z11Factor to +1.9. Nope... this did not work. Put configuration file value of Z4Z11Factor = 0.0 and edit value in lbcfpia.pro... DID NOT WORK EITHER. Edited lbcfpia_read_cfg.pro to set Z4Z11Factor = +1.9. But this did not work either... How can we change the initial z4/z11 scaling? One more place where it can be changed: sendsph.pro... edited this file to change -1.9 to +1.9. This did the trick!

09:26 Set Matrix Element 12 to -4 (and restored lbcfpia_read_cfg.pro lbcfpia.pro and lbcfpia.cfg to original settings).

09:28 Run dofpia with changes mentioned above (09:26).

10:11 (Missing some notes from the last 30 minutes)

A recap: We edited the sendsph.pro file to change z4/z11 scaling from +1.9 to 0.9 and then -1.0 (where it is now). Blue Matrix element 12 was changed from -2.4 to -4.

At 09:55 we had a problem with the Blue rotator. Error messages are:

2011/02/26 10:08:54.185902 N B ROTATOR           GOYA bad communication echo "$" instead of " ": RETRY
2011/02/26 10:08:54.286000 N B CAMERA   CCDCTRL  not available electronics temperature reading
2011/02/26 10:08:54.305905 W B ROTATOR           GOYA bad communication echo "$" instead of " " [src/rotator/rotator.c:2025]
2011/02/26 10:08:54.305939 W B ROTATOR           GOYA bad communication with serial device [src/rotator/rotator.c:1520]
2011/02/26 10:08:54.305999 N B ROTATOR           moving from -61.776 deg to -52.523 deg position
2011/02/26 10:08:54.425940 N B ROTATOR           GOYA bad communication echo "$" instead of " ": RETRY
2011/02/26 10:08:54.545936 W B ROTATOR           GOYA bad communication echo "$" instead of " " [src/rotator/rotator.c:2025]

10:13 Restarted LBC Blue. Run dofpia.

10:37 Had to close for wind. May reopen and continue later as this tuning of the Z4/Z11 interdependence for LBC Blue was not completed. Two values we had been tweaking are:(after a new set of values, exit and restart IDL)
  • in lbcfpia.cfg, the 12 element of the matrix, set to -4 now: line 112: BlueZIMatrix = [[1.0,-4.0,-8.0],[0.0,1.0,3.0],[0.0,0.0,1.0]]
  • in sendsph3.pro: the factor that multiplies res[10], -1.0 now: line 23: res[3] = -1.0 * res[10] ; 20110226

10:49 We need to fix the BlueZ22Zero still. On a stable night we are forcing in a large amount of Z22 - we can see this on the plot of forces.

11:29 Wind gusts up to 27m/s. Storm is approaching us. We call it a night.

-- MarcoPedani - 25 Feb 2011
Topic revision: r9 - 01 Mar 2011, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback